Delph Company uses a job-order costing system and has two manufacturing departments-Molding and Fabrication. The company provRequired 2A Required 2B Required 2C Required 2D Compute the departmental predetermined overhead rates. (Round the final answeRequired 2A Required 2B Required 2C Required 2D Compute the total manufacturing cost assigned to Job D-70 and Job C-200. (RouRequired 2A Required 2B Required 2C Required 2D ........ .. If Delph establishes bid prices that are 150% of total manufacturRequired 2A Required 2B Required 2C Required 2D What is Delphs cost of goods sold for the year? (Round your intermediate cal
Answer #1

2a) Overhead rate :

Molding department = (790000/28000+6) = 34.21 per MH

Fabrication department = (270000/38000+6) = 13.11 per MH

2b) Total manufacturing cost

Job D-70 Job C-200
Direct material 697000 540000
Direct labor 340000 380000
Molding overhead 684200 273680
Fabrication overhead 104880 393300
Total 1826080 1586980

2c) Bid price

Job D-70 = 1826080*1.5 = 2739120

Job C-200 = 1586980*1.5 = 2380470

2d) Cost of goods sold = 1826080+1586980 = 3413060